OpenStudy (anonymous):

Jen is 13 years younger than Andre. The sum of their ages in years is 137. What is Jen’s age in years?

OpenStudy (anonymous):

So we have A+J=137

OpenStudy (anonymous):

We need to solve for one of the terms because we are missing both. We have J in terms of A so let's re-write J. J=A-13

OpenStudy (anonymous):

now substitute

OpenStudy (anonymous):

A+A-13=137

OpenStudy (anonymous):

which we can write as 2A-13=137

OpenStudy (anonymous):

add 13 to both sides and we get 2A=150

OpenStudy (anonymous):

divide both sides by 2 and we get A=75

OpenStudy (anonymous):

now we substitute for A back in our original problem (A+J=137)

OpenStudy (anonymous):

and we get 75+J=137

OpenStudy (anonymous):

62

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

so her age is 62?

OpenStudy (anonymous):

Sure is

OpenStudy (anonymous):

I thought she would be younger. Ha. Thanks a bunch!

OpenStudy (anonymous):

Yea most problems don't go with people that old. It was a little odd